solve using INTERGRATION BY PARTS technique
c. To =
Answer #1

c) s tona dx By integration by parts - Sudv=ur-Svau o Let uz [hnx] a v= 1 / dx . By differentiating I By integrating . Sarasu=2 v7 dox-2 ()+c bt / =25x lnx-2 (2) x2 +C = 2 x dn (x) -45x4c 1 ... lnx ax=252 un xl-45a+C) csscanned with CamScanner